OpenStudy (nino_jim):

a man runs at a speed of 6.5 mi/hr. how long does it take him to run 6000 meters?(1 mile =m1690 meters)

OpenStudy (revircs):

You may want to post this to the Physics part of OpenStudy, but Ill help anyways. So we have a speed of 6.5mi/hr. Were going to want to convert this into m/hr since 1mi=1690m, the speed it 10985m/hr from there, we can set up a proportion. 10985m/1hr =6000/x then solve for x to get the time in hours.